OpenStudy (anonymous):

@StudyGurl14 So surface area again in terms of pi.

OpenStudy (studygurl14):

Okay remember. Formula for surface area of cylinder: Lateral Area + area of bases \(\large 2\pi rh+2r^2\pi\)

OpenStudy (studygurl14):

Diameter = 2 So...radius = 1 height = 9

OpenStudy (anonymous):

I'm getting 2 * pi * 2 * 9 + 2 * 4 pi right?

OpenStudy (anonymous):

where did you get the one from?

OpenStudy (anonymous):

right you simplify.

OpenStudy (anonymous):

ohhh okay. Yeah

OpenStudy (studygurl14):

hold on, it would actually be... \(\large 2\pi(1)(9)+2\pi(1)^2\)

OpenStudy (studygurl14):

because the radius is 1, not 2

OpenStudy (anonymous):

okay right

OpenStudy (studygurl14):

now... \(\large 2(9)\pi+2(1)\pi\rightarrow 18\pi+2\pi= ???\)

OpenStudy (anonymous):

Do I multiply 18 by 2 or add them?

OpenStudy (anonymous):

add then.

OpenStudy (studygurl14):

add

OpenStudy (anonymous):

alright 20 pi.

OpenStudy (studygurl14):

yep

OpenStudy (studygurl14):

i have to go now and eat lunch. Can you handle the rest?

OpenStudy (anonymous):

You are the best!! I wish I could repay you for putting up with me and all my questions. Thank you!

OpenStudy (anonymous):

Yes thank you!

OpenStudy (studygurl14):

:)